Please answer all parts

ing from nanocoulombs to microcoulombs. k here for iled view m Status ▲ 50% Part (a) How many electrons are needed to form a charge of Q/--10 nC? sin0 cotan asin co 4 5 6 atan) acotan0 sinh0 Degrees ORadians I give up! Hints: 0 for a-deduction. Hints remaining: 50% Part (b) How many electrons must be removed from a neutral object to leave a net c
50% Part (a) How many electrons are needed to form a charge of Q1-10 nC? 50% Part (b) How many electrons must be removed from a neutral object to leave a net charge of Q2 1.1 C? sin0 cotan0 asin0 acos0 atan0 acotan0 sinh0 cosh tanh0 cotanh0 O Radians VO Submit I give up! Hints for a 0% deduction. Hints remaining:- Feedback: 1% deduction per feedback -
Answer #1

from quantization of charge

n = q / e

a)

n = 10*10^-9 / ( 1.6*10^-19)

n = 6.25*10^10 electrons

b)

n = 1.1*10^-6 / ( 1.6*10^-19)

n = 6.875*10^12 electrons

========

Comment in case any doubt, will reply for sure.. Goodluck